# Single-shot and measurement-based quantum error correction via fault complexes

### v1

@@ Appendix C Single-shot decoding for fault complexes
$D_{X}$ $=\left(\begin{array}[]{ccccc}{\color[rgb]{0,.5,.5}M_{X}}&&&&\\ {\color[rgb]{0,.5,.5}\mathbbm{1}}&{\color[rgb]{0,.5,.5}H_{X}}&{\color[rgb]{0,.5,.5}\mathbbm{1}}\\ &&{\color[rgb]{0,.5,.5}M_{X}}&&\\ &&\mathbbm{1}&{\color[rgb]{0.68,0,0}H_{X}}&{\color[rgb]{0.68,0,0}\mathbbm{1}}\\ &&&&{\color[rgb]{0.68,0,0}M_{X}}\end{array}\right).$ (37)
In the above matrix, the odd rows detect ancilla faults and the even rows detect data faults, while odd columns are ancilla fault locations and even columns are data fault locations. Unlike in the CBQC framework for single-shot error correction, in the fault complex picture, there is no distinction between data faults and ancilla faults, or equivalently between detectors arising from stabilizers ($H_{X}$) and detectors arising from meta-checks ($M_{X}$). Thus, for any error $e_{X}$ with syndrome $s_{X}=D_{X}e_{X}$ we have $\partial_{1}s_{X}=\partial_{1}\partial_{2}e_{X}=0$.
In Eq. 37 we have also highlighted two blocks. The first block (teal) is the relevant check matrix for a single round of the overlapping window decoder, specifying a single detector block in the overlapping window decoder for $w=1$. The second block (red) is identical to the single-stage decoding matrix proposed in Ref. [22]. This block can be recovered from block one by treating the ancilla faults from the previous round (first column of block one) as perfect, which eliminates them from the decoding problem. The full block is the relevant detector check matrix for a window size $w=2$.
Appendix D Numerical simulations
@@
Figure 4: Example of the threshold fitting procedure for the 3D toric code under the photonic GKP noise model with an $w=2$ overlapping window decoder after $129$ rounds of noisy syndrome measurements. Left: Histogram of the bootstrap resamples, with the mean and $99\%$ confidence interval highlighted in blue. Center: Collapse plot of the fitted data. Right: Threshold plot of the data obtained from the sustainable threshold simulation with the threshold value highlighted as a black dashed line and $99\%$ confidence intervals as a grey shading.
D.1 Phenomenological noise
@@
We begin by describing the case of phenomenological noise. Since we are only considering a single type of Pauli error in our simulations, we only need to decode a single side of the fault complex. First, we construct the associated fault complex through foliation based on the base code, e.g., the 3D toric code, over $\delta$ rounds. In the cluster state picture, this corresponds to a cluster state of $12\delta+$ 1 layers. As we are using 30 iterations of the min-sum belief propagation (BP) algorithm, potentially followed by ordered statistics decoding post-processing with an order-60 combination sweep (CS-60) search strategy, we do not require the gauge group generators. Given the dual detectors $D_{Z}$, we sample faults corresponding to each matrix column from an i.i.d noise model with probability $p$. Any faults occurring on the initial and final layer are removed to assert that we are starting and ending in a code state. This assumption is equivalent to the usual assumption in the phenomenological noise model simulations that the final syndrome measurement is noise-free. For each trial, the sampled fault vector is sliced up according to the $(w,1)$-overlapping window decoder before being passed to the local decoder. A more explicit description of this can be found in Refs [48, 50]. After the full correction $c_{\delta}$ over $\delta$ rounds is calculated, we check whether the residual error is a logical error. For the 3D and 4D toric codes that possess 3 and 6 logical qubits, we count as a failure if at least a single logical qubit remains in error.
D.2 Photonic GKP noise
@@
When simulating Xanadu’s photonic architecture based on optical Gottesman-Kitaev-Preskill qubits [1, 2, 4], a few adjustments to the simulation routine are necessary. First, as we are aiming to simulate an approximation of the physical noise model, we now require knowledge about the bi-adjacency matrix of the (qubit-level) fault-tolerant graph state. As described in Ref. [4], using the process of macronization and stiching the qubit-level graph state can equivalently be described by several GKP Bell pairs and linear-optical circuit elements, e.g., phase shifters and beam splitters. For each optical mode describing a GKP state, we sample a random displacement from the normal distribution $\mathcal{N}(0,\sigma^{2})$ with variance $\sigma^{2}$. This is an effective model to describe a finite-energy GKP qunaught state with $-10\log_{10}(\sigma^{2}/\sigma_{\text{vac}}^{2})\;\mathrm{dB}$ of per-peak squeezing, where we set the vaccum variance $\sigma_{\text{vac}}^{2}=2$. As described in more detail in Refs. [1, 2, 4], the sampled displacements can be processed to yield measurement outcomes for the individual modes when they are measured in the reduction step described in Ref. [4]. From this, an effective qubit-level noise model can be obtained and the rest of the simulation routine continues as described in the previous section.
D.3 Threshold estimation
@@
$g(x)=a\left[1-\left(1-\frac{1+\tanh(bx)}{2}\right)^{c}\right],$ (40)
@@
where $x$ is the rescaled error rate as in Eq. (38) and $a,b,c$ are free fit parameters. In principle, if the fit is performed over a sufficient range of $x$, one does not need to fit $a$ and can set it instead to its analytical value $a\equiv 1-(0.5)^{k}$; however, we leave $a$ as a free parameter. The parameter $c$ typically scales with the number of decoding rounds.
@@
To estimate the threshold based on the model Eq. (40), we use the bootstrap resampling technique with $10000$ resamples; see [75] for an implementation. The obtained (asymmetric) error bars represent the $99\%$ confidence intervals. Fig. 4 contains some example plots for the threshold fitting procedure.
D.4 Stability experiments
In the 3D toric code stability experiments described in the End Matter of the main text, we observe the expected behaviour for a fixed code and increasing window size. Namely, we observe exponential suppression of the logical error rate below threshold; see Fig. 5.
@@ D.4 Stability experiments
Figure 5: Stability experiments with the $L=5$ 3D toric code. (a) Logical error rate as a function of the Pauli error rate $p$ for various sizes of the decoding window $w$. (b) Scaling of the logical error rate below the threshold as a function of the size of the decoding window $w$.
Appendix E Fault complexes of subsystem codes
@@
Prior work on foliation [9, 77, 10] gives us clues for the correct form of the fault complexes of subsystem codes. We know that in a foliated CSS subsystem code the ancilla qubit connections correspond to the rows of $G_{X}$ and $G_{Z}$ and the data qubits connections are unchanged. This suggests that—for a length-3 fault complex—the boundary operator $\partial_{2}$ describing the cluster state connectivity should be as in Eq. 31 with $\partial_{i}^{C}=G_{Z}^{T}$ and $\partial_{i-1}^{C}=G_{X}$. However, for the other boundary operators, we should use $H_{X}$ and $H_{Z}$ instead of $G_{X}$ and $G_{Z}$, as subsystem code detectors are formed from the stabilizers not the checks. Explicitly,
$D_{X}=\partial_{1}$ $=\left(\begin{array}[]{c|c}\mathbbm{1}_{r}\otimes H_{X}&R\otimes\mathbbm{1}_{n_{0}}\end{array}\right),$ (42)
$\partial_{2}$ $=\left(\begin{array}[]{c|c}\mathbbm{1}_{r}\otimes G_{Z}^{T}&R\otimes\mathbbm{1}_{n_{1}}\\ \hline\cr 0&\mathbbm{1}_{c}\otimes G_{X}\end{array}\right),$ (45)
$D_{Z}^{T}=\partial_{3}$ $=\left(\begin{array}[]{c}R\otimes\mathbbm{1}_{n_{2}}\\ \hline\cr\mathbbm{1}_{c}\otimes H_{Z}^{T}\end{array}\right).$ (48)
For the case of CSS subsystem codes with meta-checks, we expect that $D_{X}$ and $D_{Z}$ would simply need to include $M_{X}$ and $M_{Z}$ (the meta-check matrices) and would resemble Eq. 37. Given the similarity of the detector matrices, the decoding problem for these codes should be identical in structure to the decoding problem for fault complexes we have previously covered. It would be therefore interesting to examine the performance of, e.g., subsystem codes with single-shot error correction [36, 45, 68, 78] using the approach explored in this work.
The changes outlined above are relatively minor and we therefore anticipate foliated CSS subsystem codes can be represented naturally as fault complexes. The main barrier to formalizing our intuition is finding the appropriate representation of CSS subsystem codes as chain complexes (see [79, 68] for some possibilities). We leave this for future work.
Appendix F Fault complexes of circuit models
@@
$G_{X}$ $=\left(\begin{array}[]{cc}\mathbf{a}_{X}\otimes\mathbbm{1}_{n}&1_{m_{X}^{C}}\otimes\mathbf{H}_{X}^{T}\\ \mathbf{d}_{X}^{T}\otimes H_{Z}&0\end{array}\right),$ (51)
$G_{Z}$ $=\left(\begin{array}[]{cc}\mathbf{a}_{Z}\otimes\mathbbm{1}_{n}&1_{m_{Z}^{C}}\otimes\mathbf{H}_{Z}^{T}\\ \mathbf{d}_{Z}^{T}\otimes H_{X}&0\end{array}\right),$ (54)
@@
$D_{X}$ $=\left(\mathbbm{1}_{m_{X}^{B}}\otimes H_{X}\quad\mathbf{a}_{X}^{T}\otimes\mathbbm{1}_{r_{X}}\right),$ (55)
$D_{Z}$ $=\left(\mathbbm{1}_{m_{Z}^{B}}\otimes H_{Z}\quad\mathbf{a}_{Z}^{T}\otimes\mathbbm{1}_{r_{Z}}\right),$ (56)
@@
$L_{X}=(\mathbf{g}_{X}\otimes\ell_{X}),$ (57)
$L_{Z}=(\mathbf{g}_{Z}\otimes\ell_{Z}),$ (58)
@@
More evidence that our formalism can also describe fault-tolerant error correction is found by considering the explicit example of the $r$-times repeated measurement of stabilizer generators. This example is discussed in Sec. VII C of Ref. [32]. For this example, translating the notation of Li to our notation, $\mathbf{a}_{X}$ is the check matrix of the repetition code and $g_{X}$ is the all-one vector, equivalent to what we have obtained in Section B from the Künneth formula. The matrix $\mathbf{d}_{X}$ is equivalent to the $r$-dimensional identity matrix $\mathbbm{1}_{r}$ up to an additional all-zero row. Similar expressions are obtained when expressing $D_{Z}$ and $L_{Z}$.

### v2

@@ Acknowledgements.
$D_{X}$ $=\left(\begin{array}[]{ccccc}{\color[rgb]{0,.5,.5}M_{X}}&&&&\\ {\color[rgb]{0,.5,.5}\mathbbm{1}}&{\color[rgb]{0,.5,.5}H_{X}}&{\color[rgb]{0,.5,.5}\mathbbm{1}}\\ &&{\color[rgb]{0,.5,.5}M_{X}}&&\\ &&\mathbbm{1}&{\color[rgb]{0.68,0,0}H_{X}}&{\color[rgb]{0.68,0,0}\mathbbm{1}}\\ &&&&{\color[rgb]{0.68,0,0}M_{X}}\end{array}\right).$ (S25)
In the above matrix, the odd rows detect ancilla faults and the even rows detect data faults, while odd columns are ancilla fault locations and even columns are data fault locations. Unlike in the CBQC framework for single-shot error correction, in the fault complex picture, there is no distinction between data faults and ancilla faults, or equivalently between detectors arising from stabilizers ($H_{X}$) and detectors arising from meta-checks ($M_{X}$). Thus, for any error $e_{X}$ with syndrome $s_{X}=D_{X}e_{X}$ we have $\partial_{1}s_{X}=\partial_{1}\partial_{2}e_{X}=0$. This distinction runs deep in the CBQC community and has historically influenced decoder design, particularly in phenomenological noise simulations where faults are sampled in separate steps. In such simulations, faults on data qubits are first drawn from the space $C_{i}$ , and only afterwards are ancilla qubit faults sampled from $C_{i\pm 1}$. This sequential treatment reinforces the mindset that data and ancilla faults are fundamentally distinct and should be handled separately. Consequently, decoder designs—especially for single-shot codes with metachecks—have traditionally relied on heuristics to merge these distinct fault types into an effective decoding matrix-[23, 24]. By contrast, the fault complex framework eliminates this conceptual divide. By embedding data and measurement faults in the same vector spaces, $F_{i}$ and $F_{i+1}$ for primal and dual faults, respectively, it provides a natural and systematic way to construct the appropriate detector check matrices. This is illustrated inEq. S25 where the structure of $D_{X}$ naturally recovers the effective parity-check matrix for decoding. Rather than relying on intuition, as was done in prior works, the fault complex formalism directly specifies the correct boundary operators, yielding the appropriate check structure from first principles. The first block (teal) is the relevant check matrix for a single round of the overlapping window decoder, specifying a single detector block in the overlapping window decoder for $w=1$. The second block (red) is identical to the single-stage decoding matrix proposed in Ref. [24]. This block can be recovered from block one by treating the ancilla faults from the previous round (first column of block one) as perfect, which eliminates them from the decoding problem. The full block is the relevant detector check matrix for a window size $w=2$.
Thus, the fault complex formalism provides a rigorous derivation of the single-shot decoding strategy—resolving prior uncertainties about how to construct the appropriate parity-check matrix for decoding single-shot codes with metachecks.
S2.1 Insights on decoding
Formalizing the simulation framework in the language of fault complexes yielded important insights in the improved decoding of single-shot code with metachecks. In the following, we will explain based on the example of 4D toric code how those insights were obtained from analyzing the fault complex of the code.
The 4D toric code of linear lattice size $L$ has code distance $d_{C}=L^{2}$ and metacheck distance $d_{M}=L$. A code with distance $d$ can correct up errors up to weight $\lfloor(d-1)/2\rfloor$. Then, over a window size $w\in\mathbb{N}$ the distance of the corresponding fault complex $d(\mathcal{F})$ according to Eq. (S14) is given by $d(\mathcal{F})=\min[wL,L^{2}]$. Then, for $w=1$ and $L$ odd, one would not expect $L+1$ to have improved error suppression over $L$ as the weight of correctable errors is identical. However, here and in Ref. [24] numerical results have shown that $L+1$ does indeed improve performance, which appears paradoxical from this perspective.
Our formalism allowed us to understand this apparent paradox by recognizing that the distance $\min[wL,L^{2}]$ has contributions from logical correlations of different roles. First, the emphtime-like distance, which scales as $wL$, and the space-like distance, which scales as $L^{2}$.
In Ref. [24] as well as in Figure 2 of the main text, memory experiments are performed which only space-like logical errors cause logical failures. Nevertheless, also in memory experiments, time-like errors can occur and accumulate over repeated rounds of error correction but remain unproblematic as long as they are not converted into space-like errors. However, the decoder can convert accumulated time-like errors into space-like errors. This conversion process becomes significant when analyzing single-shot thresholds, necessitating simulations over many rounds of error correction to observe threshold convergence.
We find that convergence of threshold estimates is much slower for $w=1$, where the time-like distance is minimal. Due to the increased time-like distance for $w>1$, when a time-like fault is converted into a space-like fault, this fault is more likely to directly cause a logical failure. This is the intuitive explanation for the observation of the faster convergence of the threshold. In the case $w=L$, the time-like and space-like distances become identical, and convergence is immediate and it is not necessary to simulate multiple error correction rounds in principle. However, due to boundary effects, even in that case convergence should be checked carefully.
S3 Numerical simulations
@@
S3.1 Phenomenological noise
@@
We begin by describing the case of phenomenological noise. Since we are only considering a single type of Pauli error in our simulations, we only need to decode a single side of the fault complex. First, we construct the associated fault complex through foliation based on the base code, e.g., the 3D toric code, over $\delta$ rounds. In the cluster state picture, this corresponds to a cluster state of $2\delta+1$ layers. As we are using 30 iterations of the min-sum belief propagation (BP) algorithm, potentially followed by ordered statistics decoding post-processing with an order-60 combination sweep (CS-60) search strategy, we do not require the gauge group generators. Given the dual detectors $D_{Z}$, we sample faults corresponding to each matrix column from an i.i.d noise model with probability $p$. Any faults occurring on the initial and final layer are removed to assert that we are starting and ending in a code state. This assumption is equivalent to the usual assumption in the phenomenological noise model simulations that the final syndrome measurement is noise-free. For each trial, the sampled fault vector is sliced up according to the $(w,1)$-overlapping window decoder before being passed to the local decoder. A more explicit description of this can be found in Refs [54, 56]. After the full correction $c_{\delta}$ over $\delta$ rounds is calculated, we check whether the residual error is a logical error. For the 3D and 4D toric codes that possess 3 and 6 logical qubits, we count as a failure if at least a single logical qubit remains in error.
S3.2 Photonic GKP noise
Additionally, we investigated the threshold under a noise model for Xanadu’s photonic architecture based on Gottesman-Kitaev-Preskill (GKP) qubits [75, 1, 2], with cluster state construction and error model details provided below and in [5, 6]. For the 4D toric code, we observe a threshold of approximately $10.35$ dB, comparable to that of the 2D variant [1, 2]. The 3D code achieves a threshold of approximately $7.95$ dB. In both cases, the $w=3$ decoder archives thresholds comparable to the $w=L$ decoder, see Fig. S1(a) and Fig. S1(b).
Figure S1: Sustainable thresholds for 3D (a) and 4D (b) toric codes under a noise model for a photonic GKP-based architecture. Markers represent the threshold for a $(w,1)$-overlapping window decoder as a function of noisy syndrome rounds. The black dashed line shows the threshold for the optimal window choice $w=L$. In the 3D case, we only consider primal faults, whereas in the 4D case, we consider both primal and dual faults.
Figure S2: Example of the threshold fitting procedure for the 3D toric code under the photonic GKP noise model with an $w=2$ overlapping window decoder after $129$ rounds of noisy syndrome measurements. Left: Histogram of the bootstrap resamples, with the mean and $99\%$ confidence interval highlighted in blue. Center: Collapse plot of the fitted data. Right: Threshold plot of the data obtained from the sustainable threshold simulation with the threshold value highlighted as a black dashed line and $99\%$ confidence intervals as a grey shading.
@@
When simulating Xanadu’s photonic architecture based on optical Gottesman-Kitaev-Preskill qubits [1, 2, 5, 6], a few adjustments to the simulation routine are necessary. First, as we are aiming to simulate an approximation of the physical noise model, we now require knowledge about the bi-adjacency matrix of the (qubit-level) fault-tolerant graph state. As described in Ref. [5, 6], using the process of macronization and stiching the qubit-level graph state can equivalently be described by several GKP Bell pairs and linear-optical circuit elements, e.g., phase shifters and beam splitters. For each optical mode describing a GKP state, we sample a random displacement from the normal distribution $\mathcal{N}(0,\sigma^{2})$ with variance $\sigma^{2}$ and propagate the sampled displacement through the linear optical circuit until the final homodyne measurement stage. Computationally, this propagation of sampled displacements amounts to a series of matrix vector multiplications that introduce local correlations between the modes. This is an effective model to describe a finite-energy GKP qunaught state with $10\log_{10}(\sigma^{2}/\sigma_{\text{vac}}^{2})\;\mathrm{dB}$ of per-peak squeezing, where we set the vaccum variance $\sigma_{\text{vac}}^{2}=2$. As described in more detail in Refs. [1, 2, 5, 6], the sampled displacements can be processed to yield measurement outcomes for the individual modes when they are measured in the reduction step described in Ref. [5]. To improve the logical error suppression and the observed threshold, we employ as an inner decoder on the level of the GKP code an algorithm different from the binning strategy mentioned in Refs. [1, 2]. Our correlation-aware inner decoder, described in detail in Ref. [6, see the supplementary information Section VI.A], reduces the effective error probability of individual qubits in the cluster state. From this, an effective qubit-level noise model can be obtained and the rest of the simulation routine continues as described in the previous section.
S3.3 Threshold estimation
@@
$g(x)=a\left[1-\left(1-\frac{1+\tanh(bx)}{2}\right)^{c}\right],$ (S28)
@@
where $x$ is the rescaled error rate as in Eq. (S26) and $a,b,c$ are free fit parameters. In principle, if the fit is performed over a sufficient range of $x$, one does not need to fit $a$ and can set it instead to its analytical value $a\equiv 1-(0.5)^{k}$; however, we leave $a$ as a free parameter. The parameter $c$ typically scales with the number of decoding rounds.
@@
To estimate the threshold based on the model Eq. (S28), we use the bootstrap resampling technique with $10000$ resamples; see [86] for an implementation. The obtained (asymmetric) error bars represent the $99\%$ confidence intervals. Fig. S2 contains some example plots for the threshold fitting procedure.
S3.4 Stability experiments
To further support our claim that 3D and 4D toric codes are compatible with single-shot lattice surgery, we perform simulations of stability experiments. Stability experiments can be used as a proxy for simulating lattice surgery because they probe the accuracy of reliably inferring the outcomes of an extensive number of stabilizers [25]. We perform numerical simulations over $32$ decoding rounds for a 3D toric code of linear size $L$ (with periodic boundaries in time to avoid finite-size effects). In this case, the corresponding fault complex is equivalent to the chain complex that describes an asymmetric 4D toric code with dimensions $L\times L\times L\times w$ for each decoding window. Decoding proceeds with an $(w,1)$ overlapping window decoder. In this case, the effective primal fault distance is $d_{1}=wL$, i.e., for $w\leq L$ the distance is limited by the size of the decoding window. But even for $w=1$ we expect to observe an error threshold based on the distance scaling $d\sim L$. The $w=1$ case corresponds to single-shot lattice surgery, but we note that any $w=O(1)$ corresponds to constant-time overhead lattice surgery. Our results are summarized in Fig. S3. As expected from the distance calculation, we observe an error threshold for all values of $w$, with the logical error rate decreasing with increasing $L$ below the threshold. By increasing the decoding window size $w$ from $w=1$ in Fig. S3(a), to larger, but constant, values of $w=2$ in Fig. S3(b) and $w=3$ in Fig. S3(c) we observe increased error thresholds in the stability experiments. We note that choosing $w$ adaptively on the code size, e.g., $w=L$, balances time-like and space-like errors. In practice, this allows one to find a trade-off between the number of syndrome measurement rounds required for lattice surgery, and thus the time overhead for logical gates, and the achievable logical error rates of the gate. We emphasize that these observations are in stark contrast to stability experiments with non-single-shot codes, as performed by Gidney [25] on the 2D surface code, where the effective distance is independent of $L$. For two codes with $L_{2}>L_{1}$, the code with $L_{1}$ will always have a lower error rate than the code with $L_{2}$ if decoded over the same window size $w$.
Figure S3: Logical error rate for the time-like logical correlation of a 3D toric code for different window sizes of the overlapping window decoder. In (a) $w=1$, (b) $w=2$, and (c) $w=3$. Vertical black dashed lines indicate the fitted threshold. The shading indicates a confidence interval, indicating hypotheses whose likelihoods are within a factor of 1000 of the maximum likelihood estimate.
Additionally, as expected, we observe exponential suppression of the logical error rate below threshold; see Fig. S4.
@@
Figure S4: Stability experiments with the $L=5$ 3D toric code. (a) Logical error rate as a function of the Pauli error rate $p$ for various sizes of the decoding window $w$. (b) Scaling of the logical error rate below the threshold as a function of the size of the decoding window $w$.
In the main text, we also proposed a joint memory and stability experiment for the 2D toric code with periodic boundary conditions (this would also work in higher dimensions). This experiment proceeds in the same way as a memory experiment, namely
  1. 1.
    Initialize all qubits in the $|0\rangle$ state.
  2. 2.
    Measure the stabilizers $T$ times.
  3. 3.
    Measure all qubits in the $Z$ basis.
After decoding, we can then estimate the final values of the logical $Z$ operators, and the product of all the $X$ stabilizers (in a single round). These observables correspond, respectively, to the logical correlations of the fault complex descending from the logical $Z$ operators of the 2D toric code (memory), and to the logical correlation descending from the redundancy in $H_{X}$ (stability). Each of these observables should be deterministic in the absence of errors, so using this experiment, we can estimate the probability of a spacelike logical error that disrupts the memory observable or a timelike logical error that disrupts the stability observable. We can vary the code distance $d$ and the number of rounds $T$ in order to estimate the threshold for both types of observable.
S4 Fault complexes of subsystem codes
@@
Prior work on foliation [11, 82, 12] gives us clues for the correct form of the fault complexes of subsystem codes. We know that in a foliated CSS subsystem code, the ancilla qubit connections correspond to the rows of $G_{X}$ and $G_{Z}$, and the data qubits connections are unchanged. This suggests that—for a length-3 fault complex—the boundary operator $\partial_{2}$ describing the cluster state connectivity should be as in Eq. S19 with $\partial_{i}^{C}=G_{Z}^{T}$ and $\partial_{i-1}^{C}=G_{X}$. However, for the other boundary operators, we should use $H_{X}$ and $H_{Z}$ instead of $G_{X}$ and $G_{Z}$, as subsystem code detectors are formed from the stabilizers, not the checks. Explicitly,
$D_{X}=\partial_{1}$ $=\left(\begin{array}[]{c|c}\mathbbm{1}_{r}\otimes H_{X}&R\otimes\mathbbm{1}_{n_{0}}\end{array}\right),$ (S30)
$\partial_{2}$ $=\left(\begin{array}[]{c|c}\mathbbm{1}_{r}\otimes G_{Z}^{T}&R\otimes\mathbbm{1}_{n_{1}}\\ \hline\cr 0&\mathbbm{1}_{c}\otimes G_{X}\end{array}\right),$ (S33)
$D_{Z}^{T}=\partial_{3}$ $=\left(\begin{array}[]{c}R\otimes\mathbbm{1}_{n_{2}}\\ \hline\cr\mathbbm{1}_{c}\otimes H_{Z}^{T}\end{array}\right).$ (S36)
For the case of CSS subsystem codes with meta-checks, we expect that $D_{X}$ and $D_{Z}$ would simply need to include $M_{X}$ and $M_{Z}$ (the meta-check matrices) and would resemble Eq. S25. Given the similarity of the detector matrices, the decoding problem for these codes should be identical in structure to the decoding problem for fault complexes we have previously covered. It would be, therefore, interesting to examine the performance of, e.g., subsystem codes with single-shot error correction [42, 51, 73, 79] using the approach explored in this work.
The changes outlined above are relatively minor, and we therefore anticipate that foliated CSS subsystem codes can be represented naturally as fault complexes. The main barrier to formalizing our intuition is finding the appropriate representation of CSS subsystem codes as chain complexes (see [80, 73] for some possibilities). We leave this for future work.
S5 Fault complexes of circuit models
@@
$G_{X}$ $=\left(\begin{array}[]{cc}\mathbf{a}_{X}\otimes\mathbbm{1}_{n}&1_{m_{X}^{C}}\otimes\mathbf{H}_{X}^{T}\\ \mathbf{d}_{X}^{T}\otimes H_{Z}&0\end{array}\right),$ (S39)
$G_{Z}$ $=\left(\begin{array}[]{cc}\mathbf{a}_{Z}\otimes\mathbbm{1}_{n}&1_{m_{Z}^{C}}\otimes\mathbf{H}_{Z}^{T}\\ \mathbf{d}_{Z}^{T}\otimes H_{X}&0\end{array}\right),$ (S42)
@@
$D_{X}$ $=\left(\mathbbm{1}_{m_{X}^{B}}\otimes H_{X}\quad\mathbf{a}_{X}^{T}\otimes\mathbbm{1}_{r_{X}}\right),$ (S43)
$D_{Z}$ $=\left(\mathbbm{1}_{m_{Z}^{B}}\otimes H_{Z}\quad\mathbf{a}_{Z}^{T}\otimes\mathbbm{1}_{r_{Z}}\right),$ (S44)
@@
$L_{X}=(\mathbf{g}_{X}\otimes\ell_{X}),$ (S45)
$L_{Z}=(\mathbf{g}_{Z}\otimes\ell_{Z}),$ (S46)
@@
More evidence that our formalism can also describe fault-tolerant error correction is found by considering the explicit example of the $r$-times repeated measurement of stabilizer generators. This example is discussed in Sec. VII C of Ref. [36]. For this example, translating the notation of Li to our notation, $\mathbf{a}_{X}$ is the check matrix of the repetition code and $g_{X}$ is the all-one vector, equivalent to what we have obtained in Section S1 from the Künneth formula. The matrix $\mathbf{d}_{X}$ is equivalent to the $r$-dimensional identity matrix $\mathbbm{1}_{r}$ up to an additional all-zero row. Similar expressions are obtained when expressing $D_{Z}$ and $L_{Z}$.
S6 “Foliation” beyond the repetition code
We consider as an example a classical linear code that extends the repetition code. In particular, we focus on a code family defined by the polynomial $t(x)=1+x+x^{2}$. Note that, in a similar notation, the ordinary repetition code is defined by the polynomial $1+x$.
The polynomial $t(x)$ defines a family of classical linear codes with parameters $[n=3s,k=2,d=2s]$, where the parameter $s>0$ determines the size $3s\times 3s$ of the circulant matrices used in the code construction.
Codewords of the classical code take the form
$\ell_{1}$ $=\underbrace{(1,\dots,1)}_{s\text{ times}}\otimes(1,1,0),$ (S47)
$\ell_{2}$ $=\underbrace{(1,\dots,1)}_{s\text{ times}}\otimes(1,0,1).$ (S48)
Compared to the ordinary repetition code, this code encodes two logical bits, at the cost of a reduced distance.
We compute the (co-)homology of the chain complex whose boundary map is associated with the check matrix $\tilde{\mathcal{R}}$ of this code. The resulting homology groups are
[…]
As a result, the number of logical correlations and logical errors in the fault complex is increased. To be explicit, assuming for simplicity that the base CSS code has full-rank check matrices, the primal logical errors $H_{i}(\mathcal{F})$ and the primal logical correlations $H^{i}(\mathcal{F})$ are given by
$\begin{split}H_{i}(\mathcal{F})&\cong H_{0}(\tilde{\mathcal{R}})\otimes H_{i}(\mathcal{C})\oplus H_{1}(\tilde{\mathcal{R}})\otimes H_{i-1}(\mathcal{C})\\ &=\langle((1,0,\ldots,0)\otimes\ell_{Z},(0,1,0,\ldots,0)\otimes\ell_{Z},\mathbf{0})\rangle,\end{split}$ (S53)
where $\ell_{Z}\in\ker(\partial_{i})/\mathrm{im}(\partial_{i+1})$.
Similarly, the primal logical correlations are
$\begin{split}H^{i}(\mathcal{F})&\cong H^{0}(\tilde{\mathcal{R}})\otimes H^{i}(\mathcal{C})\oplus H^{1}(\tilde{\mathcal{R}})\otimes H^{i-1}(\mathcal{C})\\ &=\langle(\ell_{1}\otimes\ell_{X},\ell_{2}\otimes\ell_{X},\mathbf{0})\rangle,\end{split}$ (S54)
where $\ell_{X}\in\ker(\partial_{i+1}^{T})/\mathrm{im}(\partial_{i}^{T})$.
We emphasize that, due to the structure of the check matrix $\tilde{R}$—as evident from its polynomial description $t(x)=1+x+x^{2}$—each bit is involved in three checks. As a consequence, the space-time code described by $\mathcal{F}$ is non-matchable, even if the underlying (static) CSS code chain complex $\mathcal{C}$ is associated with a matchable code.
Additionally, we note that, while the graph state / fault complex now hosts an additional logical correlation and logical error, it is not immediately obvious how to operate on this correlation. A similar issue arises in constructing CSS codes with multiple logical qubits.
Nevertheless, it may be beneficial to consider codes beyond the repetition code for “foliation” when the base code is already a non-matchable LDPC code. In such cases, the decoding problem already requires algorithms such as BP-OSD, making this extension a natural choice.